OpenStudy (anonymous):

Find all solutions to the equation. cos^2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

clueless?? Help!!!

jimthompson5910 (jim_thompson5910):

let z = cos(x), so the equation turns into z^2 + 2z + 1 = 0 if you solve z^2 + 2z + 1 = 0, what solutions do you get?

OpenStudy (anonymous):

well it looks like the basic form to the quadratic formula, do i need to use the quadratic formula?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes that is a good method to use

jimthompson5910 (jim_thompson5910):

you can also factor, but factoring doesn't always work which is why the quadratic formula is a better method.

OpenStudy (anonymous):

Here is one

OpenStudy (anonymous):

thank you @Allenismyname and @jim_thompson5910 i figured it out the trig fuctions confused me